A 0.015-kg marble moving to the right at 0.40 m/s has a head-on, elastic collision with a 0.045 kg marble sitting at rest on a smooth level surface. Which of the following are the correct magnitudes and directions of the velocities of the two marbles after collision?
1
Expert's answer
2020-11-10T06:58:55-0500

From the Law of Conservation of Momentum, we have:


"m_1v_{1i}+m_2v_{2i}=m_1v_{1f}+m_2v_{2f},"

here, "m_1=0.015\ kg" is the mass of the first marble, "m_2=0.045\ kg" is the mass of the second marble, "v_{1i}=0.40\ \dfrac{m}{s}" is the initial velocity of the first marble before the collision, "v_{2i}=0\ \dfrac{m}{s}" is the initial velocity of the second marble before the collision, "v_{1f}" is the final velocity of the first marble after the collision, "v_{2f}" is the final velocity of the second marble after the collision.

Since collision is elastic, kinetic energy is conserved and we can write:


"\dfrac{1}{2}m_1v_{1i}^2+\dfrac{1}{2}m_2v_{2i}^2=\dfrac{1}{2}m_1v_{1f}^2+\dfrac{1}{2}m_2v_{2f}^2."

This formula gives us an additional relationship between velocities. Therefore, with the help of these two formulas we can find the velocity of each marbles after collision:


"v_{1f}=\dfrac{(m_1-m_2)}{(m_1+m2)}v_1,""v_{1f}=\dfrac{(0.015\ kg-0.045\ kg)}{(0.015\ kg+0.045\ kg)}\cdot 0.40\ \dfrac{m}{s}=-0.20\ \dfrac{m}{s}."

The sign minus means that the first marble moves to the left at "0.20\ \dfrac{m}{s}."


"v_{2f}=\dfrac{2m_1v_1}{(m_1+m_2)},""v_{2f}=\dfrac{2\cdot 0.015\ kg\cdot 0.40\ \dfrac{m}{s}}{(0.015\ kg+0.045\ kg)}=0.20\ \dfrac{m}{s}."


The sign plus means that the second marble moves to the right at "0.20\ \dfrac{m}{s}."

Answer:

"v_{1f}=0.20\ \dfrac{m}{s}," left.

"v_{2f}=0.20\ \dfrac{m}{s}," right.